OpenStudy (anonymous):

The function f(t) = 4t2 - 8t + 7 shows the height from the ground f(t), in meters, of a roller coaster car at different times t. Write f(t) in the vertex form a(x - h)2 + k, where a, h, and k are integers, and interpret the vertex of f(t). f(t) = 4(t - 1)2 + 3; the minimum height of the roller coaster is 3 meters from the ground f(t) = 4(t - 1)2 + 3; the minimum height of the roller coaster is 1 meter from the ground f(t) = 4(t - 1)2 + 2; the minimum height of the roller coaster is 2 meters from the ground f(t) = 4(t - 1)2 + 2; the minimum height of the roller coaster is 1 m

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

thats what i thought, but y?

OpenStudy (anonymous):

4t2-8t+7= =4[(t-1)2-1]+7 =4(t-1)2-4+7 =4(t-1)2+3 this is a translated parabola with a=4 traslation on x=+1 trasl. on y = +3

OpenStudy (anonymous):

i should've just put it on a graph lol. thx

OpenStudy (anonymous):

yes it would have been much simpler ;)